OpenStudy (anonymous):

write a program in c++ that would accept a number and write it in fibonacci sequence

OpenStudy (ankit042):

I will help you with the logic of Fibonacci series. 0,1,1,2,3,5,8,13.... so F(n)=F(n-1)+F(n-2) You can use different methods to code this (simple for loop with two temp variable can do the trick, recursion, memorization)

OpenStudy (anonymous):

Exactly 0 + 1 = 1, 1 + 1 = 2, 1 + 2 = 3, 2 + 3 = 5, 3 + 5 = 8, 5 + 8 = 13, 8 + 13 = 21, and so on. notice the pattern?